OpenStudy (anonymous):

Sergio recorded the color of each gumball that he bought from a machine. His results are shown below. Based on Sergio's results, what is the probability of getting a red gumball from this machine? 3/10 1/3 3/7 2/3

OpenStudy (anonymous):

@kelliegirl33 and @CGGURUMANJUNATH and @perl

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

@divu.mkr

undeadknight26 (undeadknight26):

hmm...this is not a very accurate question because we do not know how much gumballs he bought.

OpenStudy (anonymous):

Sorry there's a pic

OpenStudy (anonymous):

do you know?

undeadknight26 (undeadknight26):

nope..because they didn't give me a specific color they just asked a meaningless question...who cares what color just eat them!

OpenStudy (perl):

whats the question?

OpenStudy (anonymous):

At the top read it

OpenStudy (anonymous):

Sergio recorded the color of each gumball that he bought from a machine. His results are shown below. 3/10 1/3 3/7 2/3

undeadknight26 (undeadknight26):

I don't understand the question so sorry i cannot help you...@arabpride

OpenStudy (perl):

tatum, that is not a question, that is a statement

OpenStudy (perl):

@tatumrocks do you know the difference between a question and a statement?

OpenStudy (anonymous):

The fractions are the answer choices

OpenStudy (perl):

a question usually ends with a question mark. i dont see any question mark

OpenStudy (perl):

imagine i am a robot, and i answer only questions. now give me a question.

OpenStudy (anonymous):

Sergio recorded the color of each gumball that he bought from a machine. His results are shown below. NOW LOOK AT THE CHART.

OpenStudy (anonymous):

RE-READ QUESTION.

undeadknight26 (undeadknight26):

@tatumrocks they never specifically told us the question...they just said his results are shown below...which is not a question...the graph does not answer the question: is there a question? so we cannot help you until there is a question like: what is the ratio between green gumballs to the rest.

OpenStudy (anonymous):

I said re-read the question

undeadknight26 (undeadknight26):

There is no question.

OpenStudy (perl):

tatum, you edited the question

OpenStudy (anonymous):

I know

OpenStudy (perl):

ok we want probability of red gumball. there are 12 red gumballs. probability = # favorable/ # total P(red) = 12 / ( 10 + 12 + 18) = 12 /40 = 3/10